alones (alones):

l = x + 4; w = x + 3

alones (alones):

The area of a rectangle is found by multiplying the length by the width: A = lw. A certain rectangle has an area of x 2 + 7x + 12. Factor the trinomial to find the length and width of the rectangle. my answer was the top one

OpenStudy (thephysicsman):

you can factor the polynomial because remember A = l*w so that's technically how you would go about getting it.

OpenStudy (thephysicsman):

(x+4)(x+3)

alones (alones):

x^2+7x+12

alones (alones):

Wait so that was right? l = x + 4; w = x + 3

OpenStudy (thephysicsman):

yep

alones (alones):

k thanks <3